# The Invisible Leash? Why RLVR May or May Not Escape Its Origin

### v2

@@ Conclusion.
This trade-off underpins RLVR’s empirical strengths in tasks with narrowly defined optimal solutions, such as mathematical proofs or tactical game endgames (where precision is paramount), while also emphasizing the need for explicit diversity mechanisms in more open-ended domains, such as code generation, creative writing (Feizi et al., 2023; Ding et al., 2024), or brainstorming (Chang & Li, 2025). Importantly, entropy reduction is not inherently undesirable: when a task admits a unique correct solution, lower answer-level entropy simply reflects desirable convergence. Importantly, even in multi-solution domains, concentrating mass on a narrower set may still be desirable under constrained compute budgets. However, our results show that entropy reduction can still lead to empirical support shrinkage even in predominantly single-solution domains like math, where RLVR sometimes fails to recover valid completions still accessible to the more diverse base model. This highlights that entropy-induced narrowing is a general phenomenon, not limited to multi-solution tasks, underscoring the broader need for explicit exploration or diversity-promoting strategies.

### v3

@@ Conclusion.
This trade-off underpins RLVR’s empirical strengths in tasks with narrowly defined optimal solutions, such as mathematical proofs or tactical game endgames (where precision is paramount), while also emphasizing the need for explicit diversity mechanisms in more open-ended domains, such as code generation, creative writing (Feizi et al., 2023; Ding et al., 2024), or brainstorming (Chang and Li, 2025). Importantly, entropy reduction is not inherently undesirable: when a task admits a unique correct solution, lower answer-level entropy simply reflects desirable convergence. Importantly, even in multi-solution domains, concentrating mass on a narrower set may still be desirable under constrained compute budgets. However, our results show that entropy reduction can still lead to empirical support shrinkage even in predominantly single-solution domains like math, where RLVR sometimes fails to recover valid completions still accessible to the more diverse base model. This highlights that entropy-induced narrowing is a general phenomenon, not limited to multi-solution tasks, underscoring the broader need for explicit exploration or diversity-promoting strategies.